OpenStudy (anonymous):

If h(x) = -4x - 10, find h(-5). A. -30 B. 1.25 C. 10 D. -1

OpenStudy (anonymous):

what do u think

OpenStudy (anonymous):

B? 1.25?

OpenStudy (anonymous):

one method is process of elimation

OpenStudy (solomonzelman):

So I have: \(\color{#000000 }{ \displaystyle h(x)=-4x-10 }\) What does that tell me? This function tells me, that for any number "x" that I input into the function, to product the output, I will multiply that number times -4 and subtract ten from the result. So, for example if I need to find the output when x=4, (or f(4)), then I would plug in 4 directly into the function h(x), instead of x. \(\color{#000000 }{ \displaystyle h(2)=-4\cdot 2-10=-8-10=-18 }\)

OpenStudy (anonymous):

ok so use that in and equation

OpenStudy (anonymous):

Is it B?

OpenStudy (corpsss):

h(x)=-4x-10, and h(-5) let be the -5 is the x h(x)=-4(-5)-10 h(x)=20-10 h(x)=10 so the answer is letter C. 10